Fermi National Laboratory


Accelerator Update

January 5
The day shift began with store #3127 colliding, stacking in the Antiproton Source (Pbar), and with beam to MiniBoone.

Linac RF station #3 tripped off at 7:53 PM and would not reset, an expert was called in to investigate. The station was back on at 9:13 PM.

Switchyard 120 (SY120) studies began at 9:25 PM and ended at 1:17 AM.

Tuesday January 6
The day shift began with store #3127 colliding, stacking, and with beam to MiniBoone.

Store # 3127 was terminated at 7:27 AM.

Operations put the Tevatron (TeV) into shot setup at 9:11 AM. At 11:30 AM, Operations established store #3130 with an initial luminosity of 48.09 E30. Pbar started studies.

At 3:50 PM Pbar returned to stacking.

A P1 line power supply (I:HV703) tripped off at 7:12 PM and would not reset. Power supply techs were called in to work on the problem and the power supply was back on at 9:13 PM.

Wednesday January 7
The Midnight shift began with store #3130 colliding, stacking, Recycler Ring (RR) studies, and with beam to MiniBoone.

At 1:05 AM a Booster Ion Pump tripped off causing some of the Booster RF stations to trip off. All station were back on by 1:28 AM.

Today's Plans
End the store this evening, followed by shot setup, then return to stacking.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/7/2004   email Fermilab